Lead is a poisonous metal that especially affects children because they retain a larger fraction of lead than adults do. Lead levels of 0.250ppm in a child cause delayed cognitive development. How many moles of lead present in 1.00g of child’s blood would 0.250ppm represent.


1
Expert's answer
2021-09-21T07:51:16-0400

"nPb = \frac{mPb}{mmPb} = \frac{ppm(Pb)×msolution}{10^6ppm×mm(Pb)}"


"= \frac{0.250ppm×1.00g}{10^6×207.2g/mol}"


= 1.21mol